OpenStudy (anonymous):

What is the value of the discriminant in the equation shown below? x^2-4x+10=0

OpenStudy (anonymous):

The discriminant is b^2 - 4ac. In your proble, a = 1, b = - 4, and c = 10. Therefore, b^2 - 4ac = (-4)^2 - 4(1)(10) = 16 - 40 = - 36

OpenStudy (***[isuru]***):

if the quadratic function is ax^2 + bx + c = 0 then the discriminant is b^2 - 4ac so... in ur question x^2 - 4x +10=0 ax^2 + bx + c = 0 discriminant is... b^2 - 4ac (-4)^2 - 4(10) = 16 - 40 = -36

OpenStudy (anonymous):

ok awesome i have one more question....

OpenStudy (anonymous):

Describe the roots of the equation shown below. 3x^2+4x+2=0